ADNAN MUHAMMED

Grade 12,

Length of a rectangle is increased by 50% and breadth is decreased by 25% what is the difference in the area

Length of a rectangle is increased by 50% and breadth is decreased by 25% what is the difference in the area

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Let length =l

breadth=b

New length=1.5l

new breadth=0.75b

New area 1.5l*0.75b

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sudheesh Singanamalla
13 years ago

let length be l and breadth be b

area = l*b

new length = l + l/2 = 3l /2

new breadth = b - b/4 = 3b/4

new area = 3l/2 * 3b/4 = 9 * l * b / 8 ;

 

difference in area = 9*l*b/8 - l*b

==> (l*b) * (9/8 - 1)

==> (l*b) * 1/8

thefore difference in area = 1/8 * l*b

 

please approve if the answer is correct